### v2

# Pareto-Grid-Guided Large Language Models for Fast and High-Quality Heuristics Design in Multi-Objective Combinatorial Optimization

Abstract
Multi-objective combinatorial optimization problems (MOCOP) frequently arise in practical applications that require the simultaneous optimization of conflicting objectives. Although traditional evolutionary algorithms can be effective, they typically depend on domain knowledge and repeated parameter tuning, limiting flexibility when applied to unseen MOCOP instances. Recently, integration of Large Language Models (LLMs) into evolutionary computation has opened new avenues for automatic heuristic generation, using their advanced language understanding and code synthesis capabilities. Nevertheless, most existing approaches predominantly focus on single-objective tasks, often neglecting key considerations such as runtime efficiency and heuristic diversity in multi-objective settings. To bridge this gap, we introduce MPaGE, a novel enhancement of the Simple Evolutionary Multiobjective Optimization (SEMO) framework that leverages LLMs and Pareto Front Grid (PFG) technique. By partitioning the objective space into grids and retaining top-performing candidates to guide heuristic generation, MPaGE utilizes LLMs to prioritize heuristics with semantically distinct logical structures during variation, thus promoting diversity and mitigating redundancy within the population. Through extensive evaluations, MPaGE demonstrates superior performance over existing LLM-based frameworks, and achieves competitive results to traditional Multi-objective evolutionary algorithms (MOEAs), with significantly faster runtime. Our code is available at https://github.com/langkhachhoha/MPaGE.
1 Introduction
Multi-objective combinatorial optimization problems (MOCOP)commonly arise in real-world applications such as vehicle routing, production planning, where multiple conflicting objectives must be optimized simultaneously over a large, discrete solution space (Türkyılmaz et al. 2020; Liu et al. 2020; Phan Duc et al. 2025). Unlike single-objective optimization, MOCOP aims to approximate the Pareto front, which captures trade-offs among non-dominated solutions.
Due to the NP-hard nature of these problems, exact algorithms are often impractical, resulting in the widespread use of heuristic and metaheuristic methods such as NSGA-II, MOEA/D, and MOEA/D-DE (Deb et al. 2002; Zhang and Li 2007a; Li and Zhang 2008). Despite their success, these methods typically rely on domain-specific knowledge and extensive iterative search. Several neural approaches have been proposed for MOCOP, aiming to automatically learn heuristics and improve adaptability (Zhang et al. 2022; Hieu et al. 2024; Fan et al. 2024). However, these methods often require retraining for different problem sizes, demand substantial computational resources, and struggle to generalize to problems with unseen input formats.
Recently, Large Language Models (LLMs)have shown strong capabilities in automatic heuristic design, offering a new paradigm for optimization algorithm development (Wu et al. 2024; Liu et al. 2025; Novikov et al. 2025). By leveraging their language understanding and code generation abilities, LLMs can produce heuristics and the corresponding implementations with minimal human intervention. Recent approaches integrate LLMs with evolutionary strategies to iteratively evolve effective problem-solving programs (Liu et al. 2024; Ye et al. 2024; Romera-Paredes et al. 2024a; van Stein and Bäck 2024). While achieving competitive performance and often surpassing traditional methods, most existing LLM-based evolution frameworks primarily target single-objective problems, with limited exploration of their applicability to multi-objective settings. The inherent complexity and trade-off structure of MOCOP introduce significant challenges that require tailored and robust design strategies. To address this, Huang et al. (Huang, Zhang, and Liu 2025) propose an LLM-based framework for discovering and refining evolutionary operators suited to diverse MOCOP. However, runtime efficiency, which is crucial for the real-world deployment of MOCOP solvers, remains an underexplored aspect of LLM-driven heuristic design. MEoH (Yao, Chen, and Wang 2025) considers both optimality and efficiency within a multi-objective evolutionary framework. Additionally, prior LLM-based methods tend to produce populations of algorithms with similar operational logic and slight representation difference, limiting both population diversity and the creativity of LLMs.
To address the aforementioned issues, we propose MPaGE, a novel framework designed to solve MOCOP while simultaneously discovering a pareto front of LLM-generated heuristics that jointly optimize solution quality and runtime efficiency, with an explicit emphasis on promoting heuristic diversity. Our approach curates heuristic algorithms for the Simple Evolutionary Multiobjective Optimization (SEMO)paradigm, leveraging Pareto Front Grid (PFG)to guide the design of LLM-based variation heuristics by partitioning the objective space into grids, retaining leading individuals from promising regions, and enhancing both solution quality and search efficiency. From these regions, MPaGE builds a pool of elitist candidates and employs LLMs to assess their semantic structures, clustering them into groups of similar logic. Variation is then performed with respect to these clusters, promoting semantic diversity and mitigating redundancy within the heuristic population. To the best of our knowledge, this is the first comprehensive evaluation of LLM-generated heuristics on standard MOCOP, addressing both solution quality, computational efficiency and semantic diversity. Our main contributions are as follows:
  • We introduce MPaGE, the first framework to systematically combine LLMs with the SEMO paradigm and PFG, aiming to solve MOCOP by jointly optimizing runtime, solution quality, and maintaining semantic diversity.
  • We leverage LLMs to verify the logical structure of heuristics and perform cross-cluster recombination, thereby enhancing diversity and reducing redundancy through logically dissimilar variations.
  • We conduct extensive experiments on standard MOCOP benchmarks, demonstrating consistent improvements in runtime efficiency, solution quality, and semantic diversity over LLM-based baselines and traditional MOEAs.
2 Related Works
2.1 Multi-objective Optimization Algorithms
Solving MOCOP often involves extending metaheuristics to handle multiple objectives in discrete domains. NSGA-II (Deb et al. 2002) and MOEA/D (Zhang and Li 2007a) remain popular for effectively maintaining Pareto front diversity. Local search techniques like Pareto Local Search (PLS) (Paquete and Stützle 2004), SEMO (Laumanns, Thiele, and Zitzler 2004) improve exploitation by iteratively exploring non-dominated neighbors. More recent neural approaches aim to learn the entire Pareto set, as in PMOCO (Lin, Yang, and Zhang 2022), or boost diversity through dual mechanisms, as in NHDE (Chen et al. 2023b). Despite their success in solving MOCOP, these methods still rely on handcrafted components, struggle to generalize across diverse problem settings, and incur substantial computational costs.
2.2 LLMs for Heuristic Design
LLMs are increasingly used to automate heuristic design in combinatorial optimization. Core approaches like EoH, AEL, and FunSearch employ evolutionary frameworks to generate, combine, and refine heuristics in natural language or code, outperforming traditional methods on TSP, bin packing, and scheduling tasks (Liu et al. 2024, 2023; Romera-Paredes et al. 2024b). Similarly, HSEvo incorporates harmony search principles into the LLM-driven generation process to promote diversity and adaptability in solutions (Dat, Doan, and Binh 2025). Other work combines LLMs with synthesis techniques like MCTS for guided search (Zheng et al. 2025). While promising, most of these methods primarily focus on single-objective tasks, with limited attention to MOCOP, neglects other practical criteria like efficiency and complexity (Gutjahr 2012).
2.3 Multi-Objective Optimization with LLMs
Recent LLM-based heuristics have been applied to meta-heuristic design. Huang et al. (Huang, Zhang, and Liu 2025) used LLMs to generate crossover and mutation operators for multi-objective optimization, later extending this to evolutionary multitasking with LLM-designed knowledge transfer models (Huang et al. 2024). Nevertheless, these methods overlook the computational cost of generating and evaluating heuristics, a key challenge in MOEAs. MEoH, REMoH (Yao, Chen, and Wang 2025; Forniés-Tabuenca et al. 2025) present a multi-objective evolutionary framework that optimizes multiple performance metrics, such as optimality and efficiency, to discover trade-off algorithms in a single run. However, they struggle to distinguish heuristics with similar logic but different implementations, reducing diversity on the pareto front and hindering multi-objective exploration.
3 Preliminary
3.1 Multi-Objective Combinatorial Optimization
Generally, we consider MOCOP defined as: $\min_{x\in\mathcal{X}}\;f(x)=(f_{1}(x),\ldots,f_{M}(x)),$ where $\mathcal{X}$ is a finite or discrete feasible set, and $f\colon\mathcal{X}\to\mathbb{R}^{M}$ maps each solution to $M$ objectives to be minimized. Trade-offs among objectives are characterized by Pareto optimality, as defined below, and solution quality is commonly assessed using the hypervolume (HV) indicator (Zitzler and Thiele 1999).
Definition 1 (Pareto dominance). For solutions $x^{a},x^{b}\in\mathcal{X}$, $x^{a}$ dominates $x^{b}$, denoted $x^{a}\prec x^{b}$, if $f_{i}(x^{a})\leq f_{i}(x^{b})$ for all $i\in\{1,\ldots,M\}$, and there exists some $j$ such that $f_{j}(x^{a})<f_{j}(x^{b})$.
Definition 2 (Pareto optimality). A solution $x^{*}\in\mathcal{X}$ is Pareto optimal if no $x\in\mathcal{X}$ satisfies $x\prec x^{*}$. The set of all such solutions forms the Pareto set ($\mathrm{PS}$), and Pareto front $\mathrm{PF}=\{{f}(x)\mid x\in\mathrm{PS}\}$.
Figure 1: Overview of the proposed method. The heuristics are first represented in a feature space, and then partitioned into grid cells based on their objective values. Potential parent heuristics are then chosen with the assistance of LLM-based review from grids. Crossover and mutation operations are applied, guided by informative feedback from LLM-based reflection to enhance population diversity. Finally, non-dominated individuals are retained to form the next generation for the subsequent iteration.
3.2 Simple Evolutionary Multiobjective Optimization
Simple Evolutionary Multiobjective Optimization (SEMO)is a minimalist yet effective MOEAs, widely used in theoretical analyses of runtime performance (Laumanns, Thiele, and Zitzler 2004). Similar to Pareto Local Search (PLS) (Paquete and Stützle 2004), SEMO maintains an archive of nondominated solutions. As shown in Algorithm 1, each iteration selects a random solution from the archive, explores one of its neighbors, and updates the archive if a new nondominated solution is discovered. Unlike PLS, which evaluates all neighbors, SEMO samples only one neighbor per iteration, which may limit its anytime performance. See appendix B for detailed insights compared to other frameworks.
$s\leftarrow\textit{rand\_generation}()$ ;
  // Initialize randomly a solution
$A\leftarrow\{s\}$;
repeat
 $s\leftarrow\textbf{{selection}}(A)$ ;
    // Random $s$ from AA
 $s^{\prime}\leftarrow\textbf{{neighborhood\_exploration}}(s)$ ;
    // Select a neighbor of $s$ randomly
 if $\nexists a\in A:a\prec s^{\prime}$ then // Check dominance
    $A\leftarrow A\cup\{s^{\prime}\}\setminus\{a\in A\mid s^{\prime}\prec a\}$;
until stop_condition()stop\_condition();
return AA;
Algorithm 1 Simple Evolutionary Multiobjective Optimization
4 Methodology
4.1 Problem Formulation
We formalize heuristics design for MOCOP as a Language Multi-Criteria Heuristic Design (LMHD) problem, aiming to discover heuristics that balance evaluation criteria reflecting various aspects of heuristic behavior, such as effectiveness, efficiency, or generalization across problems. The key components are introduced as follows:
Multi-Criteria Evaluation Function. We define $E:\mathcal{H}\rightarrow\mathbb{R}^{M}$ to evaluate each heuristic $h\in\mathcal{H}$ over $M$ criteria, capturing its expected performance and key behavioral aspects on a given MOCOP.
Language Multi-Criteria Heuristic Design. Language Multi-Criteria Heuristic Design (LMHD) is a variant of hyper-heuristic that leverages LLMs to generate a diverse set of heuristics. The goal is to discover heuristics that balance multiple criteria simultaneously. Given the heuristic space $\mathcal{H}$, the evaluation function $E(h)=\left(e_{1}(h),e_{2}(h),\ldots,e_{M}(h)\right)$ assigns to each heuristic $h$ a vector of $M$ criteria to be minimized. The aim is to approximate the Pareto front of nondominated heuristics in $\mathcal{H}$.
Our goal is to design MOCOP-specific heuristics for the selection and neighborhood exploration steps, corresponding to lines 4 and 5 in Algorithm 1. The proposed LMHD takes problem specifications as input and generates heuristics optimized over two practical and complementary criteria: solution quality and computational efficiency. Specifically, we define:
$E:\mathcal{H}\rightarrow\mathbb{R}^{2},\quad E(h)=\left(e_{1}(h),e_{2}(h)\right),$ (1)
where $e_{1}$ is the average negative hypervolume across a set of instances, measuring solution quality, and $e_{2}$ is the total running time. These two criteria reflect critical trade-offs in real-world applications, where high performance must be balanced with efficiency. Each heuristic $h\in\mathcal{H}$ is a complete algorithm that, given the current population $A$, returns a new candidate solution $s^{\prime}$ by internally performing both selection and neighborhood generation, as in Algorithm 1:
$h:2^{\mathcal{|S|}}\rightarrow\mathcal{S},\quad s^{\prime}=h(A),$ (2)
where $\mathcal{S}$ denotes the space of feasible solutions.
4.2 Overview
MPaGE integrates LLMs into a multi-phase evolutionary framework to design diverse and effective heuristics, as illustrated in Figure 1. The process begins by initializing a population of heuristics tailored for SEMO paradigm. Each heuristic is represented in a feature space defined by solution quality and running time, and the population is progressively refined until the stopping criterion is met, yielding a set of non-dominated heuristics. At each iteration, PFG generates grids that partition the objective space, thereby grouping heuristics into distinct regions (Section 4.3). A pool of elitism heuristics is selected from these grids, and an LLM-based review analyzes the logical semantics of candidate heuristics to cluster them into behaviorally similar groups (Section 4.4). Subsequently, search operators such as crossover and mutation are applied to these clusters, guided by informative feedback from LLM-based reflection, to generate new offspring. These offspring are added to the population, and non-dominated individuals are selected to form the next generation for the subsequent iteration (Section 4.5).
4.3 Pareto Front Grid mechanism
Dominance-based approaches like NSGA-II (Deb et al. 2002) and decomposition-based methods such as MOEA/D (Zhang and Li 2007b) often struggle to balance convergence and diversity in tracking the true Pareto Fronts, particularly in heuristic design scenarios where objectives like runtime and solution quality exhibit irregular or non-uniform distributions. To address these limitations, we adopt the Pareto Front Grid (PFG) approach (Xu et al. 2023a) to achieve a better balance between convergence and diversity. Guiding the search using leading solutions in PFG helps focus on promising regions, improving solution quality and efficiency while reducing redundancy.
PFG Generation: Motivated by (Xu et al. 2023a), let the objective space be denoted by $\mathcal{Z}\subset\mathbb{R}^{2}$. At generation $t$, let $\mathcal{P}^{(t)}=\{{E}({h}_{i})\in\mathbb{R}^{2}\mid{h}_{i}\in\mathcal{H}^{(t)}\}$ be the set of objective vectors of the current population, where each ${h}_{i}$ is a heuristic algorithm and $\mathcal{H}^{(t)}$ denotes the heuristic space. We define the ideal and nadir points for each objective $j\in\{1,2\}$ as $z^{*}_{j}=\min_{h\in\mathcal{P}^{(t)}}e_{j}(h)$ and $z^{n}_{j}=\max_{h\in\mathcal{P}^{(t)}}e_{j}(h)$, where $e_{j}(h)$ denotes the $j$-th objective value of heuristic $h$. The objective space is scaled via min-max normalization:
$\tilde{{E}}=\left(\frac{e_{1}-z^{*}_{1}}{z^{n}_{1}-z^{*}_{1}},\frac{e_{2}-z^{*}_{2}}{z^{n}_{2}-z^{*}_{2}}\right)^{\top}.$ (3)
To structure the population in the normalized objective space, we partition $0,1]^{2}$ into a grid of cells of side length $\delta_{1}>0$, $\delta_{2}>0$ along each objective axis, respectively, depending on the number of segments along each dimension. For each solution ${h}_{i}$, assign its objective vector $\tilde{{E}}({h}_{i})$ to a grid cell $G({h}_{i})\in\mathbb{N}^{2}$, and define the PFG mapping as:
$G({h}_{i})=\left(\left\lfloor\frac{e_{1}({h}_{i})}{\delta_{1}}\right\rfloor,\ \left\lfloor\frac{e_{2}({h}_{i})}{\delta_{2}}\right\rfloor\right)\in\mathbb{N}^{2}.$ (4)
This discretization forms a structured grid over the objective space, grouping solutions into distinct regions. We define the following mapping:
$\mathcal{G}:\mathbb{N}^{2}\to 2^{\mathcal{|H|}},\quad\mathcal{G}(g)=\left\{{h}_{i}\in\mathcal{H}^{(t)}\mid G({h}_{i})=g\right\}.$ (5)
where each cell $g$ contains all solutions whose objective vectors fall within that grid region. For each non-empty cell $g\in\operatorname{dom}(\mathcal{G})$, retain a representative subset $\mathcal{R}_{g}\subset\mathcal{G}(g)$, selected using non-dominated sorting within the cell. The union of all representatives yields the elite set:
$\mathcal{E}^{(t)}=\bigcup_{g}\mathcal{R}_{g}.$ (6)
PFG Selection: To facilitate reproduction, the population is organized into a grid, where heuristics exhibiting similar running time or solution quality are placed in neighboring cells. This spatial organization encourages crossover between heuristics with related characteristics, promoting the inheritance and refinement of useful traits.
Selection for mating is performed over a pool $\mathcal{P}$ formed using the grid structure of the objective space. With probability $\epsilon$, a group of elitism candidates is selected from a set of adjacent grid cells. Specifically, a grid cell $g$ is randomly sampled, and the pool is formed as the union of $\mathcal{G}(g)$ and its neighboring cells $\mathcal{G}(g^{\prime})$, where $g^{\prime}$ denotes the cells adjacent to $g$ along both objective axes, as shown in Figure 2. Otherwise, the parents are selected from entire $\mathcal{E}^{(t)}$. Formally:
$\mathcal{P}=\begin{cases}\mathcal{G}(g)\cup\mathcal{G}(g^{\prime}),&\text{if }\mathcal{U}[0,1]<\epsilon,\\ \mathcal{E}^{(t)},&\text{otherwise},\end{cases}$ (7)
This hybrid strategy balances local exploitation and global exploration, promoting diversity while guiding search toward the Pareto front. See Appendix C for details.
4.4 Semantic clustering
Previous methods struggle to maintain effective diversity in multi-objective heuristic design, especially for complex MOCOP. The standard MOEAs relies on Pareto dominance in the objective space $f(x)$, which is unreliable due to the stochastic nature of the heuristic performance. MEoH addresses this by introducing a dominance-dissimilarity measure based on Abstract Syntax Trees (ASTs) (Yao, Chen, and Wang 2025; Neamtiu, Foster, and Hicks 2005). However, ASTs fail to capture semantic similarity between heuristics with similar logic but different implementations, resulting in redundant individuals and reduced diversity.
Figure 2: LLM-based review clustering applies elitism heuristics based on semantic similarity and employs a probability matrix to guide variation among clusters.
To address these challenges, instead of relying solely on syntactic features, we query LLMs to assess the semantic and behavioral similarity among elite heuristics $P=\{h_{1},h_{2},\ldots,h_{n}\}$, and group them into coherent clusters as depicted in Figure 2:
$\text{SemClust}(P)$ $=\{C_{1},C_{2},\ldots,C_{m}\}$ (8)
s.t. $\bigcup\nolimits_{i=1}^{m}C_{i}=P,$
$C_{i}\cap C_{j}=\emptyset\quad(i\neq j).$
Each cluster $C_{i}$ contains heuristics whose code segments implement similar underlying logic, despite syntactic or structural differences. This semantic clustering mitigates redundancy and enhances behavioral diversity at a higher abstraction level. During variation, we apply mutation within clusters to explore local variations, and perform crossover across clusters to combine diverse behaviors. Specifically, given a randomly selected cluster $C_{i}$ and a heuristic $h\in C_{i}$, the offspring heuristic $o$ is generated as follows:
$o\leftarrow\begin{cases}\text{Mutate}(h),&\text{if }\mathcal{U}[0,1]<\gamma,\\ \text{Crossover}(h,h^{\prime}),&\text{otherwise},\quad h^{\prime}\sim\mathcal{U}\left(\bigcup_{k\neq i}C_{k}\right),\end{cases}$ (9)
More detailed insights can be found in Appendix F.
4.5 Automatic Heuristic generation
In this section, we propose the MPaGE framework for automatic heuristic design. Examples of the prompts used at each stage are provided in the Appendix E.
Individual Representation. Following previous works (Liu et al. 2024; Ye et al. 2024), MPaGE encodes each individual as a natural language description and its corresponding code implementation in Python, both generated by LLMs, together with an associated fitness score. The fitness is evaluated over a set of problem-specific instances by measuring both performance and running time.
Population initialization. MPaGE initializes the heuristic population by querying LLMs with prompts that describe the problem and specify the signature of the heuristic function to be discovered. Specifically, each heuristic is tailored for selection mechanisms and neighborhood exploration within the SEMO paradigm, then evaluated on benchmark instances and assigned a fitness score.
Selection. MPaGE partitions the objective space into grid cells to construct an elitism pool (Section 4.3), then clusters the candidates and selects parent pairs based on dissimilarity as detailed in Section 4.4, or sample from the entire population as Equation 7.
Feedback reflection. Feedback reflection enhances heuristic design by providing LLMs with clear, interpretable signals that guide improvement. For each pair of heuristic parents, an LLM-based reflection module analyzes their respective strengths and weaknesses, returning a textual suggestion on how to improve or effectively combine them.
Crossover and Mutation. MPaGE prompts the generator LLM to create an offspring heuristic by recombining or modifying parent elements, guided by reflective feedback. The prompt includes: (i) task specifications, (ii) parent heuristics in code and natural language, (iii) reflection guidance, and (iv) instructions to generate the new heuristic.
Population Management. Offspring are merged into the population, with non-dominated individuals forming the next generation, as defined by Equation 6. The process repeats until the maximum number of iterations is reached.
5 Experiments
5.1 Experimental Setup
Benchmarks.
We evaluate the proposed MPaGE framework on four widely recognized MOCOP that are extensively investigated in the literature: Bi-TSP, Tri-TSP, Bi-CVRP, and Bi-KP. Comprehensive descriptions of these problems are provided in Appendix A.
Experiment settings.
The generated heuristics are evaluated on 10 instances for each problem, with sizes of 20, 20, 50, and 50 for Bi-TSP, Tri-TSP, Bi-CVRP and Bi-KP, following the setup in (Chen et al. 2023a). The experimental parameters are configured as follows: the number of generations is set to 20, and the population size is 10 for all problems. Each crossover operator selects 2 parent heuristics to generate offspring heuristics. All heuristics are designed and evaluated under the SEMO search paradigm, with a runtime constraint of 2000 iterations and a time limit of 60 seconds. The probabilities $\epsilon=0.9$, $\gamma=0.3$, and the number of PFG segments is set to 4. Experiments were conducted on a Mac with an Apple M1 chip and 8 GB RAM. GPT-4o-mini (temperature 0.7) was used as the pre-trained LLM, while GPT-4o was used for assessing and clustering task.
Performance Metric.
Objectives
To evaluate a heuristic, we consider two objectives, both averaged across instances and minimized simultaneously: 1) Negative Hypervolume (NHV): Minimizing the negative hypervolume guides the search toward high-quality solutions. 2) Running Time: Measures the execution time of the heuristic to encourage efficiency.
Metrics
Solution quality is evaluated using the hypervolume (HV) and Inverted Generational Distance (IGD). HV indicates how well the Pareto front is approximated (higher is better), while IGD measures proximity and distribution relative to a reference front. Specifically, when comparing LLM-based methods, HV captures overall performance, as each heuristic is evaluated by NHV and execution time; for MOEAs, it reflects final solution quality. In term of diversity measurement, we use the Shannon-Wiener Diversity Index (SWDI) and Cumulative Diversity Index (CDI) (Dat, Doan, and Binh 2025). Details are provided in Appendix D.
Baseline Methods
For LLM-based automated heuristic design, we compare our method against representative approaches: EoH, MEoH, ReEvo and HSEvo (Liu et al. 2024; Yao, Chen, and Wang 2025; Ye et al. 2024; Dat, Doan, and Binh 2025) and additionally against widely adopted MOEAs, including NSGA-II, MOEA/D, SEMO and PFG-MOEA (Deb et al. 2002; Li and Zhang 2008; Xu et al. 2023b; Laumanns, Thiele, and Zitzler 2004). See Appendix B for the full experimental setup and descriptions.
Figure 3a: Pareto Front of heuristics on Bi-TSP20
Figure 3b: HV and IGD comparison on Bi-TSP20
Figure 3c: Pareto Front of heuristics on Tri-TSP20
Figure 3d: HV and IGD comparison on Tri-TSP20
Figure 3e: The non-dominated heuristics on Bi-TSP20
Figure 3f: Pareto fronts of benchmark instances, Bi-TSP 20/50
5.2 Experimental Results
Pareto Fronts and Convergence Analysis
As illustrated in Table 1, MPaGE consistently outperforms other LLM-based baselines in both convergence and Pareto front quality across all benchmarks. On Bi-TSP20 and Tri-TSP20, MPaGE achieves the highest HV scores of 0.911 and 0.936, and the lowest IGD scores of 0.010 and 0.050, respectively, indicating faster convergence and superior solution quality. Similar trends are observed on Bi-KP50 and Bi-CVRP50, where MPaGE maintains strong performance.
Table 1: Results of MPaGE compared to other baselines regarding HV and IGD on four benchmark problems. The best values are highlighted in bold.
Method Bi-TSP20 Tri-TSP20 Bi-CVRP50 Bi-KP50
HV ↑ IGD ↓ HV ↑ IGD ↓ HV ↑ IGD ↓ HV ↑ IGD ↓
EoH 0.756 0.117 0.755 0.148 0.957 0.538 0.602 0.363
ReEvo 0.541 0.435 0.694 0.547 0.658 0.462 0.996 0.138
HSEvo 0.557 0.329 0.715 0.308 0.626 0.450 0.730 0.182
MEoH 0.724 0.067 0.884 0.114 0.322 0.286 0.748 0.248
MPaGE (Ours) 0.911 0.010 0.936 0.050 0.980 0.007 0.932 0.035
In Figures 3a – 3d, the non-dominated heuristics in the final population, along with the corresponding convergence curves of HV and IGD per iteration, illustrate clear differences among methods on Bi-TSP20 and Tri-TSP20, respectively. MPaGE consistently yields a broader and more diverse set of Pareto-optimal solutions, spanning wider regions of the objective space, converges faster and clearly outperforms other baselines in terms of HV and IGD. In contrast, EoH shows slower HV growth and higher IGD due to its narrow focus on minimizing performance gaps alone. MEoH, while enhancing diversity over EoH, still trails MPaGE in convergence and final quality, likely due to the complexity of MOCOP tasks and the limited impact of its diversity mechanism.
Table 2: Two top heuristics designed by MEoH and MPaGE.
Method Bi-TSP20 Tri-TSP20 Bi-CVRP50 Bi-KP50
HV ↑ Time ↓ HV ↑ Time ↓ HV ↑ Time ↓ HV ↑ Time ↓
MEoH (best) 0.603 3.265 0.410 5.790 0.241 3.383 0.344 4.139
MEoH (fast) 0.345 0.177 0.403 5.228 0.141 0.081 0.349 2.246
MPaGE (best) 0.629 4.429 0.478 8.112 0.454 0.191 0.359 2.647
MPaGE (fast) 0.451 1.304 0.411 7.590 0.151 0.063 0.354 1.367
We evaluate the two best-so-far heuristics across multiple benchmark problems (Table 2), each in two configurations: “best” (highest hypervolume) and “fast” (lowest runtime), measured at the final population. Notably, MPaGE discovers heuristics that outperform MEoH in HV. In several cases, its fast variant is quicker and more effective, highlighting MPaGE’s strength in balancing quality and efficiency.
In and out-of-distribution size generalization analysis
Table 3: Comparison results across all benchmarks against baselines. The HV and time are averaged over 50 instances.
Method Bi-TSP20 Bi-TSP50 Bi-TSP100 Tri-TSP20 Tri-TSP50 Tri-TSP100
HV $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ Speedup Time $\downarrow$
NSGA-II 0.603 1.0x 731.2 0.495 1.0x 875.5 0.419 1.0x 932.5 0.451 3.9x 252.446 0.279 1.7x 669.9 0.195 1.5x 830.3
MOEA/D 0.597 5.8x 126.644 0.473 5.5x 157.796 0.414 4.0x 240.680 0.365 4.5x 216.485 0.189 3.6x 310.583 0.123 3.1x 398.387
PFG-MOEA 0.624 1.2x 617.734 0.532 1.0x 844.914 0.457 1.0x 952.186 0.465 1.0x 984.180 0.289 1.0x 1127.46 0.222 1.0x 1229.857
SEMO 0.543 149.4x 4.894 0.284 80.5x 10.878 0.178 56.5x 16.846 0.295 143.0x 6.881 0.108 84.8x 13.290 0.065 33.2x 36.994
EoH 0.584 52.2x 14.004 0.510 18.5x 47.233 0.503 6.0x 159.156 0.420 108.7x 9.050 0.208 39.7x 28.368 0.107 29.3x 42.038
ReEvo 0.623 107.3x 6.812 0.516 86.2x 10.151 0.413 51.0x 18.653 0.427 95.0x 10.361 0.205 41.1x 27.416 0.103 14.8x 83.213
HSEvo 0.620 112.3x 6.513 0.519 95.3x 9.182 0.422 54.1x 17.590 0.420 115.6x 8.515 0.218 55.2x 20.435 0.115 20.3x 60.659
MEoH (best) 0.603 224.0x 3.265 0.480 126.2x 6.937 0.395 79.2x 12.024 0.410 170.0x 5.790 0.205 79.4x 14.197 0.111 48.8x 25.213
MPaGE (best) 0.629 165.1x 4.429 0.542 126.0x 6.950 0.442 74.9x 12.719 0.478 121.3x 8.112 0.220 76.5x 14.747 0.109 46.6x 26.394
Method Bi-KP50 Bi-KP100 Bi-KP200 Bi-CVRP20 Bi-CVRP50 Bi-CVRP100
HV $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ Speedup Time $\downarrow$ HV $\uparrow$ Speedup Time $\downarrow$
[…]
PFG-MOEA 0.358 1.9x 488.617 0.484 1.7x 596.989 0.330 2.1x 570.174 0.598 6.6x 159.867 0.435 7.0x 277.464 0.318 9.6x 359.108
SEMO 0.195 136.9x 6.766 0.144 529.4x 1.880 0.188 115.0x 10.234 0.518 8342.8x 0.126 0.205 1616.0x 1.204 0.135 944.8x 3.634
EoH 0.343 4.9x 187.455 0.463 5.1x 195.414 0.320 18.5x 63.555 0.539 6332.5x 0.166 0.443 6176.7x 0.315 0.369 5731.7x 0.599
ReEvo 0.357 368.1x 2.517 0.452 432.1x 2.303 0.202 550.2x 2.139 0.511 3822.5x 0.275 0.354 2231.3x 0.872 0.249 2257.3x 1.521
HSEvo 0.357 330.5x 2.803 0.450 318.3x 3.127 0.205 456.5x 2.578 0.515 3185.4x 0.330 0.405 1781.7x 1.092 0.305 916.0x 3.748
MEoH (best) 0.344 223.8x 4.139 0.464 219.1x 4.543 0.321 275.3x 4.275 0.503 2280.2x 0.461 0.241 575.1x 3.383 0.126 279.3x 12.291
MPaGE (best) 0.359 350.0x 2.647 0.486 535.3x 1.859 0.197 555.7x 2.118 0.568 14599.9x 0.072 0.454 10186.7x 0.191 0.422 9563.5x 0.359
We evaluate the generalization capability of MPaGE on both in-distribution and out-of-distribution larger instances of Bi-TSP (20, 50, 100, 150, 200), Tri-TSP (20, 50, 100), and Bi-KP instances (50, 100, 200), each comprising 10 instances. All hyperparameter settings are elaborated in Appendix D. As evidenced in Table 5, MPaGE consistently outperforms baseline methods across various problem settings. While competing approaches suffer significant performance degradation as problem size increases, our method maintains remarkable stability and robustness. For instance, in Bi-TSP 200, MPaGE achieves an IGD of just 0.017, substantially lower than MEoH (0.186) and EoH (0.181), indicating a much closer approximation to the true Pareto front. Similarly, in the more complex Tri-TSP 100, MPaGE attains an IGD of 0.000, whereas MEoH and EoH produce 0.101 and 0.409, respectively, further demonstrating MPaGE ’s exceptional solution quality and scalability.
Impact of Reflection and Heuristic Diversity
We conducted experiments on the Bi-TSP20 benchmark to assess the effectiveness of our LLM Reflection approach. The results, summarized in Figure 3e and Table 4(a), demonstrate the superior performance of MPaGE over baseline methods (EoH and MEoH) in both HV and IGD metrics. Remarkably, even without the reflection-based feedback, MPaGE outperforms all baselines, indicating that it inherently benefits from capturing the correlation among objectives within each grid. When enhanced with the reflection feedbacks, MPaGE exhibits further improvements, achieving the highest HV and the lowest IGD. These findings underscore the strength of our approach in promoting both convergence and diversity. To investigate the impact of our approach on population diversity, we evaluate all frameworks using the SWDI and CDI metrics, as depicted in Table 4(b). Notably, MPaGE demonstrates significantly better performance compared to the other baselines. Higher SWDI and CDI values indicate more uniform heuristic distribution and greater population diversity, promoting effective exploration. Further analysis is provided in Appendix F.
Comparison to Conventional MOEAs
We evaluate the impact of PFG on the optimization process and compare its performance with two well-established MOEAs: NSGA-II (Deb et al. 2002) and MOEA/D (Zhang and Li 2007a). As shown in Table 6, based on experiments conducted on Bi-TSP problems, MPaGE consistently outperforms the baselines in HV and IGD. These results highlight the effectiveness of the PFG mechanism. By directing the search toward the most promising regions of the objective space, the method enhances both the solution quality and the overall efficiency of the optimization process.
Table 4: (a) Effects of LLM Reflection Bi‐TSP20; (b) Comparison in Shannon–Wiener diversity index (SWDI) and cumulative diversity index (CDI).
Method HV $\uparrow$ IGD $\downarrow$
EoH 0.688 0.141
MEoH 0.659 0.122
MPaGE w/o Feedback 0.829 0.077
MPaGE 0.941 0.023
Problems Bi TSP Bi CVRP
SWDI $\uparrow$ CDI $\uparrow$ SWDI $\uparrow$ CDI $\uparrow$
EoH 0.897 1.944 1.168 2.173
ReEvo 0.647 2.133 0.943 1.908
HSEvo 0.757 1.915 1.102 1.964
MEoH 0.639 2.086 0.143 2.181
MPaGE 1.029 2.152 1.172 2.213
Table 5: Performance comparison in terms of HV and IGD on in- and out-of-distribution instances.
Problems EoH MEoH MPaGE
HV ↑ IGD ↓ HV ↑ IGD ↓ HV ↑ IGD ↓
Bi-TSP 20 0.843 0.100 0.786 0.045 0.918 0.063
Bi-TSP 50 0.293 0.154 0.450 0.113 0.972 0.020
Bi-TSP 100 0.361 0.169 0.364 0.134 0.937 0.026
Bi-TSP 150 0.351 0.216 0.364 0.159 0.985 0.026
Bi-TSP 200 0.341 0.181 0.349 0.186 0.990 0.017
Tri-TSP 20 0.755 0.148 0.884 0.114 0.936 0.050
Tri-TSP 50 0.483 0.379 0.881 0.093 0.916 0.000
Tri-TSP 100 0.447 0.409 0.810 0.101 0.897 0.000
Bi-KP 50 0.695 0.470 0.862 0.061 0.907 0.008
Bi-KP 100 0.751 0.390 0.828 0.243 0.923 0.063
Bi-KP 200 0.316 0.704 0.472 0.192 0.855 0.080
Table 6: Efficiency of our Pareto Front Grid
Backbone Bi-TSP 20 Bi-TSP 50 Bi-TSP 100
HV ↑ IGD ↓ HV ↑ IGD ↓ HV ↑ IGD ↓
NSGA-II 0.860 0.052 0.801 0.120 0.757 0.095
MOEA/D 0.819 0.119 0.768 0.108 0.560 0.157
PFG (Ours) 0.913 0.024 0.836 0.075 0.844 0.099
Evaluation Against Baselines
[…]
6 Conclusion
In this paper, we propose MPaGE, a novel LLM-guided framework for solving MOCOP that simultaneously discovers a Pareto front of heuristics balancing solution quality, runtime efficiency, and semantic diversity. Integrating LLMs with the SEMO paradigm and introducing the Pareto Front Grid, our approach efficiently partitions the objective space and steers heuristic evolution toward promising regions. By clustering heuristics based on semantic logic and promoting inter-group diversity, the framework ensures meaningful variation within the heuristic population. Empirical results show that MPaGE consistently outperforms prior LLM-based approaches in achieving superior trade-offs across objectives, enhancing heuristic diversity, and reducing computational cost. Furthermore, it outperforms traditional algorithms in efficiency while maintaining comparable solution quality, demonstrating its potential as a scalable and generalizable approach for automated heuristic discovery.
Appendix: Table of Contents
This is Appendix for “Pareto-Grid-Guided Large Language Models for Fast and High-Quality Heuristics Design in Multi-Objective Combinatorial Optimization”.
  • A  Detailed of MOCOP benchmarks . A
    • A.1  Multi-objective traveling salesman problem (MOTSP) . A.1
    • A.2  Multi-objective capacitated vehicle routing problem (MOCVRP) . A.2
    • A.3  Multi-objective knapsack problem (MOKP) . A.3
  • B  Baseline Descriptions . B
    • B.1  LLM-based Baselines . B.1
    • B.2  MOEAs Baselines . B.2
    • B.3  Motivation for Using SEMO as the Underlying Framework . B.3
  • C  Algorithm details . C
  • D  Metric descriptions . D
    • D.1  Hypervolume . D.1
    • D.2  Inverted Generational Distance . D.2
    • D.3  Shannon-Wiener diversity index . D.3
    • D.4  Cummmulative diversity index . D.4
  • E  MPaGE prompt details . E
    • E.1  Task description prompt . E.1
    • E.2  Initialization prompt . E.2
    • E.3  Semantic clustering prompt . E.3
    • E.4  Feedback reflection prompt . E.4
    • E.5  Crossover prompt . E.5
  • F  Clustering variations analysis . F
  • G  Designed Heuristics . G
  • H  Comparison to Neural Combinatorial Optimization . H
Appendix A Detailed of MOCOP benchmarks
A.1 Multi-objective traveling salesman problem (MOTSP)
In the multi-objective traveling salesman problem (MOTSP) involving $n$ nodes and $M$ objectives, each node $i\in\{1,\ldots,n\}$ is represented by $M$ sets of 2-dimensional coordinates, one for each objective. For a given objective $m$, the Euclidean distance $c_{ij}^{m}$ between two nodes $i$ and $j$ is calculated based on their corresponding coordinates. The goal is to construct a tour $\pi$ that visits every node exactly once and simultaneously minimizes the total travel distance across all $M$ objectives. Formally, the objective vector to minimize is defined as:
$f(\pi)=\left(f_{1}(\pi),f_{2}(\pi),\ldots,f_{M}(\pi)\right),$
with each component given by:
$f_{m}(\pi)=\sum_{i=1}^{n-1}c_{\pi_{i},\pi_{i+1}}^{m}+c_{\pi_{n},\pi_{1}}^{m},\quad\forall m\in\{1,\ldots,M\}.$
Problem instances are synthetically generated by sampling all coordinates independently and uniformly from the $0,1]^{2M}$ hypercube.
A.2 Multi-objective capacitated vehicle routing problem (MOCVRP)
We focus on the bi-objective capacitated vehicle routing problem (Bi-CVRP), which considers a set of $n$ customer nodes and a single depot. Each node, including the depot, is assigned a location in 2D space, and every customer has a specific demand. A fleet of identical vehicles with uniform capacity is stationed at the depot and must complete routes that collectively serve all customers, returning to the depot afterward. Each vehicle must have sufficient remaining capacity to fulfill the demand of any customer it visits. The problem simultaneously optimizes two conflicting objectives: minimizing the total traveled distance and minimizing the makespan, defined as the length of the longest route. In our Bi-CVRP instances, node coordinates are randomly sampled from the $0,1]^{2}$ space, and customer demands are randomly selected from the set $\{1,\ldots,9\}$. The vehicle capacity is set to 30, 40, and 50 for problem sizes where $20\leq n<40$, $40\leq n<70$, and $70\leq n\leq 100$, respectively. To standardize inputs, all demand values are normalized with respect to vehicle capacity.
A.3 Multi-objective knapsack problem (MOKP)
In the multi-objective knapsack problem (MOKP) with $M$ objectives and $n$ items, each item is characterized by a weight and $M$ distinct profit values, one for each objective. These items can be visualized as nodes within an instance graph. The goal is to select a subset of items such that the total profit across all $M$ objectives is maximized, while the combined weight of the selected items does not exceed a predefined capacity. Formally, let $x\in\{0,1\}^{n}$ be a binary decision vector, where $x_{i}=1$ indicates that item $i$ is selected. The problem can be stated as:
$\text{maximize }f(x)=\left(f_{1}(x),f_{2}(x),\ldots,f_{M}(x)\right),$
subject to:
$\sum_{i=1}^{n}w_{i}x_{i}\leq C,\quad x_{i}\in\{0,1\},\quad\forall i\in\{1,\ldots,n\},$
where $w_{i}$ is the weight of item $i$, and $f_{m}(x)=\sum_{i=1}^{n}p_{i}^{m}x_{i}$ represents the total profit for objective $m$, with $p_{i}^{m}$ being the profit of item $i$ under objective $m$.
The instances are generated by independently sampling all weights $w_{i}$ and profits $p_{i}^{m}$ from the uniform distribution over $[0,1]$. The knapsack capacity $C$ is set to 12.5 for $50\leq n<100$ and 25 for $100\leq n\leq 200$, respectively.
Appendix B Baseline Descriptions
B.1 LLM-based Baselines
We compare the Pareto fronts of heuristics generated by MPaGE with those of the following methods:
  • EoH (Liu et al. 2024) is a framework that combines large language models and evolutionary computation to automatically design heuristics. It simulates expert reasoning by evolving both code and thought via five tailored prompt strategies.
  • ReEvo (Ye et al. 2024) is an evolutionary framework that integrates LLMs with reflective feedback to enhance heuristic generation.
  • HSEvo (Dat, Doan, and Binh 2025) is a framework built on harmony search principles, introducing new components and enhancing evolutionary operators to jointly optimize objective performance and solution diversity.
  • MEoH (Yao, Chen, and Wang 2025) is an LLM-based framework that formulates heuristic design as a multi-objective optimization problem to generate diverse, non-dominated heuristics in a single run.
To ensure fairness, experimental parameters follow MPaGE: 20 generations and a population size of 10 across all problems. Each crossover operator selects two parent heuristics to generate offspring.
B.2 MOEAs Baselines
For all algorithms, we assume a minimisation problem as:
$\min_{x\in\mathcal{X}}\;F(x)=\bigl{(}f_{1}(x),\ldots,f_{M}(x)\bigr{)},$
where $F:\mathcal{X}\rightarrow\mathbb{R}^{M}$ is vector-valued with mutually conflicting objectives, and Pareto-dominance $\preceq$ is defined in the usual way. To benchmark the effectiveness of LLM-generated heuristics in solving multi-objective combinatorial optimization problems, we compare them against three well-established MOEAs and SEMO. These algorithms are representative of different design philosophies in evolutionary multi-objective optimization and are widely used in the literature.
NSGA-II (Deb et al. 2002)
NSGA-II maintains a population $\mathcal{P}_{t}$ of fixed size $N$. It contains components:
  • Fast non-dominated sorting. Each solution $x$ is assigned a rank
    $r(x)=\min\{k\in\mathbb{N}_{0}\mid x\in F_{k}\},$
    where $F_{k}$ is the $k$-th non-dominated front. Fronts are constructed iteratively: $F_{0}$ contains all non-dominated individuals, $F_{1}$ those dominated only by members of $F_{0}$, and so on. The total sorting cost is $O(MN^{2})$ (or $O(MN\log N)$ with incremental approaches).
  • Crowding distance. Within each front $F_{r}$, crowding distance is calculated as
    $d(x)=\sum_{m=1}^{M}\frac{f_{m}(x^{\,\text{next}})-f_{m}(x^{\,\text{prev}})}{f_{m}^{\max}-f_{m}^{\min}},$
    where $x^{\text{prev}}$ and $x^{\text{next}}$ are the nearest neighbours in front $F_{r}$ along each objective $f_{m}$, and missing values at boundaries are treated as zero.
  • Selection and variation. Binary tournament based on the lexicographic key (r(x),d(x))(r(x),-d(x)) chooses parents; simulated binary crossover (SBX) and polynomial mutation create $N$ offspring.
  • Elitist replacement. Parents and offspring are merged and the best $N$ solutions under (r,d)(r,-d) are kept, guaranteeing $O(1)$ elitism per generation.
    The algorithm explicitly balances convergence (via rank) and diversity (via distance) while preserving the worst-case $O(N)$ memory footprint.
MOEA/D (Zhang and Li 2007b)
MOEA/D decomposes the multi-objective problem into $K$ single-objective sub-problems using a set of weight vectors $\{\boldsymbol{\lambda}^{(k)}\}_{k=1}^{K}$ with $\boldsymbol{\lambda}^{(k)}\in\Delta^{M-1}$, where $\Delta^{M-1}$ is the unit simplex.
  • Scalarisation. Typical aggregation functions include:
    $g_{\text{WS}}(x\mid\boldsymbol{\lambda}^{(k)})$ $=\sum_{m=1}^{M}\lambda^{(k)}_{m}f_{m}(x),$
    $g_{\text{Tcheb}}(x\mid\boldsymbol{\lambda}^{(k)},\mathbf{z}^{\ast})$ $=\max_{m}\,\lambda^{(k)}_{m}\,\bigl{|}f_{m}(x)-z_{m}^{\ast}\bigr{|},$
[…]
    where $\mathbf{z}^{\ast}=(\min f_{1},\dots,\min f_{M})$ is the ideal point.
  • Neighbourhood collaboration. Each sub-problem $k$ is allotted a neighbourhood $\mathcal{N}_{k}$ comprising the $T$ closest weight vectors in Euclidean distance. Variation operators are restricted to parents drawn from $\mathcal{N}_{k}$, and offspring $x^{\prime}$ may update the solutions of all neighbours $j\in\mathcal{N}_{k}$ if $g(x^{\prime}\mid\boldsymbol{\lambda}^{(j)})<g(x^{(j)}\mid\boldsymbol{\lambda}^{(j)})$.
    The decomposition transforms the $M$-objective search into parallelised scalar searches of complexity $O(KT)$ per generation while implicitly promoting solution diversity through the geometry of $\{\boldsymbol{\lambda}^{(k)}\}$.
PFG-MOEA (Xu et al. 2023a)
PFG-MOEA is a recent grid-guided extension of MOEA/D that introduces PFG.
  • Pareto-front normalisation. Let $\mathbf{z}^{\ast}$ and $\mathbf{z}^{\mathrm{nad}}$ be the ideal and nadir approximations of the current population. For any solution $x$ define the normalised objective vector $\tilde{F}(x)=(f_{1}^{\prime},\dots,f_{M}^{\prime})$ with $f_{m}^{\prime}=\dfrac{f_{m}(x)-z_{m}^{\ast}}{z_{m}^{\mathrm{nad}}-z_{m}^{\ast}}$.
  • Grid construction. Given a grid size $G\in\mathbb{N}$, the cell index of $x$ is
    $\boldsymbol{g}(x)=\bigl{(}g_{1},\dots,g_{M}\bigr{)},\qquad g_{m}=\bigl{\lfloor}G\,f_{m}^{\prime}\bigr{\rfloor}.$
    Only the leading (knee) solution in each occupied cell is retained in an external PFG archive, drastically lowering memory and update costs to $O(|\text{PFG}|)$, with $|\text{PFG}|\ll N$ for moderate $G$.
  • Environmental selection. At every generation PFG-MOEA:
    1. (i)
      updates $\mathbf{z}^{\mathrm{nad}}$ via a statistical estimator using a random sample of the population, and
    2. (ii)
      performs a grid-based knee-point selection: within each $\boldsymbol{g}$ the solution maximising $\kappa(x)=\sum_{m}\bigl{|}f_{m}^{\prime}(x)-\tfrac{1}{2}\bigr{|}$ is chosen to preserve extreme trade-offs.
    The algorithm then follows MOEA/D-style neighbourhood variation and update, but the decisions are guided by the sparsified PFG archive, yielding superior convergence/diversity on irregular fronts with an empirical time complexity close to $O(KT+|\text{PFG}|)$ per generation.
SEMO (Li et al. 2024)
is a minimalistic $(1+1)$ MOEA widely used in theoretical studies. At each step, it selects a random solution from the current archive and applies a simple mutation to generate one neighbor. The new solution is added to the archive if it is not dominated, while any dominated solutions are removed. Its key strength lies in its simplicity and efficiency in exploring the Pareto front through fast, randomized local search, making it especially effective on discrete and combinatorial problems.
To ensure a fair and controlled comparison, all algorithms are configured with a population size of 300 and are evolved for 300 generations across all benchmark instances. This uniform experimental setup allows performance differences to be attributed primarily to the algorithmic design, rather than variations in computational budget or search effort. For the SEMO algorithm and the LLM-generated heuristics applied to SEMO, we run 20,000 iterations for Bi-TSP and Tri-TSP, and 10,000 iterations for Bi-KP and Bi-CVRP.
Combinatorial optimization problems are encoded appropriately based on their structure. For permutation-based problems such as Bi-TSP, Tri-TSP, and Bi-CVRP, we adopt standard permutation representations and apply combinatorial variation operators including Swap Mutation and Partially Mapped Crossover. For binary-encoded problems such as Bi-KP, we use bitstring representations and apply bit-flip mutation along with standard uniform crossover to maintain diversity and introduce variation during evolution.
B.3 Motivation for Using SEMO as the Underlying Framework
In this work, we adopt the Simple Evolutionary Multi-objective Optimization (SEMO) framework as the basis for evaluating LLM-generated heuristics, motivated by both practical efficiency and algorithmic effectiveness:
  • Lightweight and Computationally Efficient: Evaluating LLM-generated heuristics involves running optimization routines repeatedly, often hundreds or thousands of times over combinatorial problems. Embedding each heuristic within full-fledged MOEAs such as NSGA-II or MOEA/D results in significant computational overhead. In contrast, SEMO follows a $(1+1)$ design that evaluates only one solution per iteration, which greatly reduces runtime and makes large-scale heuristic search tractable.
  • Surprisingly Effective Search Performance: Despite its simplicity, SEMO has demonstrated competitive, and in some cases superior, performance compared to both traditional MOEAs and local search heuristics (Li et al. 2024). One key factor is its unbounded archive, which enables the generation of new solutions from the entire set of non-dominated individuals, in contrast to the fixed-size populations used in most MOEAs. This promotes greater diversity and adaptability in the search. Additionally, SEMO’s inherent stochasticity in selection and mutation introduces a less greedy search behavior, which helps the algorithm escape local optima more effectively than deterministic local search methods such as PLS or Anytime PLS (Dubois-Lacoste, López-Ibáñez, and Stützle 2015; Paquete and Stützle 2004).
Building on these advantages, we employ LLMs to design the selection and neighborhood exploration components within the SEMO framework, resulting in heuristics that combine speed with strong search performance for MOCOP.
Appendix C Algorithm details
In this section, we provide a detailed explanation of the PFG generation process and the proposed MPaGE Framework, as illustrated in Algorithm 2 and Algorithm 3, respectively.
1
Input: Population $\mathcal{H}$ of heuristics;
Objective values $e_{j}(h)$ for $j=1,2$;
Number of grid segments $K_{1},K_{2}$;
Small positive value $\sigma$
Output: Grid mapping $\mathcal{G}:\mathbb{N}^{2}\to 2^{\mathcal{|H|}}$; Elite set $\mathcal{E}$
2
3for $j\leftarrow 1$ to 22 do
 $z^{*}_{j}\leftarrow\min_{h\in\mathcal{H}}e_{j}(h)$ ;
    // Ideal point of objective $j$
4
 $z^{n}_{j}\leftarrow\max_{h\in\mathcal{H}}e_{j}(h)$ ;
    // Nadir point of objective $j$
5
 $\delta_{j}\leftarrow\frac{z^{n}_{j}-z^{*}_{j}+2\sigma}{K_{j}}$ ;
    // Cell width with margin
6
7
8foreach $h\in\mathcal{H}$ do
9 for $j\leftarrow 1$ to 22 do
    $g_{j}\leftarrow\left\lfloor\frac{e_{j}(h)-z^{*}_{j}+\sigma}{\delta_{j}}\right\rfloor$ ;
       // Grid index
10
11
 $G(h)\leftarrow(g_{1},g_{2})$ ;
    // Grid cell index for $h$
12
 $\mathcal{G}(G(h))\leftarrow\mathcal{G}(G(h))\cup\{h\}$ ;
    // Assign $h$ to cell
13
14
[…]
16
17foreach cell $g$ in $\mathcal{G}$ do
 $\mathcal{C}\leftarrow\mathcal{G}(g)$ ;
    // Current set of solutions in cell
18
 $\mathcal{G}(g)\leftarrow\emptyset$ ;
    // Reset cell content
19
20 foreach $h\in\mathcal{C}$ do
21    if $\nexists h^{\prime}\in\mathcal{C},\ h^{\prime}\neq h\wedge e(h^{\prime})\prec e(h)$ then
       $\mathcal{G}(g)\leftarrow\mathcal{G}(g)\cup\{h\}$ ;
          // Keep non-dominated $h$
22
23
24
$\mathcal{E}\leftarrow\bigcup_{g\in\operatorname{dom}(\mathcal{G})}\mathcal{G}(g)$ ;
  // Collect elite set from all non-empty cells
25
26return $\mathcal{G},\ \mathcal{E}$;
Algorithm 2 PFG Generation
Input: Population size NN; Iteration count $T$;
Problem description $\Pi$; Pretrained LLM $\mathcal{L}$;
Grid size $(K_{1},K_{2})$; Margin $\sigma$;
Probabilities $\epsilon$ (local selection), $\gamma$ (mutation)
Output: Final heuristic population $P^{*}$
1 $P_{0}\leftarrow\emptyset$;
2
3for $i\leftarrow 1$ to NN do
4 $o\leftarrow$ LLMGenerate($\mathcal{L},\Pi$);
5
6 $P_{0}\leftarrow P_{0}\cup\{o\}$;
7
8
9for $t\leftarrow 1$ to $T$ do
10 $\mathcal{G},\mathcal{E}\leftarrow$ PFGGeneration($P_{t-1},e_{1},e_{2},K_{1},K_{2},\sigma$);
11
12 for $i\leftarrow 1$ to NN do
13    $u\sim\mathcal{U}(0,1)$;
14
15    if $u<\epsilon$ then
       $g\sim\mathcal{U}(1,|\mathcal{G}|)$; $\mathcal{N}(g)\leftarrow$ Neighbors($g$); $\mathcal{P}\leftarrow\bigcup_{g^{\prime}\in\{g\}\cup\mathcal{N}(g)}\mathcal{G}(g^{\prime})$ ;
          // Exploration
16       $\{C_{1},\ldots,C_{m}\}\leftarrow$ SemClust($\mathcal{P};\mathcal{L}$);
17
18       $i\sim\mathcal{U}(1,m)$;  $h\sim C_{i}$;
19
20       $v\sim\mathcal{U}(0,1)$;
21
22       if $v<\gamma$ then
          $P_{\text{parent}}\leftarrow\{h\}$ ;
             // Mutation
23
24       else
25          $h^{\prime}\sim\bigcup_{k\neq i}C_{k}$;
26
          $P_{\text{parent}}\leftarrow\{h,h^{\prime}\}$ ;
             // Crossover
27
28
29
30    else
       $P_{\text{parent}}\sim\texttt{Sample}(\mathcal{E},2)$ ;
          // Exploitation
31
32
33    $\phi\leftarrow$ ReflectiveFeedback($\mathcal{L},P_{\text{parent}}$);
34
35    $o\leftarrow$ SearchOffspring($\mathcal{L},\phi$);
36
37    $P_{t-1}\leftarrow P_{t-1}\cup\{o\}$;
38
39
40
[…]
42
return $P^{*}$
Algorithm 3 MPaGE Framework
Appendix D Metric descriptions
D.1 Hypervolume
The hypervolume (HV) metric is a prevalent indicator used to assess the quality of solutions generated by multi-objective combinatorial optimization (MOCO) algorithms. It evaluates both convergence toward the Pareto front and the diversity of solutions, without relying on a known ground truth. Given a reference point $\mathbf{r}\in\mathbb{R}^{M}$, the hypervolume of a Pareto front $\mathcal{F}$ is denoted as $\mathrm{HV}_{\mathbf{r}}(\mathcal{F})$ and defined by:
$\mathrm{HV}_{\mathbf{r}}(\mathcal{F})=\mu\left(\bigcup_{\mathbf{f}(\pi)\in\mathcal{F}}[\mathbf{f}(\pi),\mathbf{r}]\right),$ (10)
where $\mu$ is the Lebesgue measure, and $\mathbf{f}(\pi),\mathbf{r}]$ represents the axis-aligned hyperrectangle spanning from the point $\mathbf{f}(\pi)$ to the reference point $\mathbf{r}$ in $M$ dimensions, i.e., $f(\pi),r]=[f_{1}(\pi),r_{1}]\times\cdots\times[f_{M}(\pi),r_{M}]$.
To ensure fair comparison of HV values across different objectives, we normalize each objective value based on global approximations. Specifically, to compare the quality of the Pareto fronts produced by LLM-based method, we compute the ideal point $\mathbf{z}^{\text{ideal}}=(z^{\text{ideal}}_{1},\ldots,z^{\text{ideal}}_{M})^{\top}$ and nadir point $\mathbf{z}^{\text{nadir}}=(z^{\text{nadir}}_{1},\ldots,z^{\text{nadir}}_{M})^{\top}$ from the union of all approximated Pareto fronts $\mathcal{P}$ obtained by all heuristics:
$f^{\prime}_{i}(\mathbf{x})=\frac{f_{i}(\mathbf{x})-z^{\text{ideal}}_{i}}{z^{\text{nadir}}_{i}-z^{\text{ideal}}_{i}},$ (11)
where $z^{\text{ideal}}_{i}=\min\{v_{i}\mid\mathbf{v}\in\mathcal{P}\},\quad z^{\text{nadir}}_{i}=\max\{v_{i}\mid\mathbf{v}\in\mathcal{P}\},\quad\forall i\in\{1,\ldots,M\}.$ This normalization maps all objective values to $[0,1]$. The HV reference point is set to $\mathbf{r}^{*}=(1.1,\ldots,1.1)^{\top}$.
To evaluate HV of a heuristic for given MOCOP, The HV is normalized as $\text{HV}^{\prime}_{\mathbf{r}}(\mathcal{F})=\text{HV}_{\mathbf{r}}(\mathcal{F})/\prod_{i=1}^{M}|r_{i}-z_{i}|$, where $\mathbf{z}$ is an ideal point such that $z_{i}<\min\{f_{i}(\pi)\mid f(\pi)\in\mathcal{F}\}$ (or $z_{i}>\max\{f_{i}(\pi)\mid f(\pi)\in\mathcal{F}\}$ for maximization), $\forall i\in\{1,\ldots,M\}$. The $\mathbf{r}$ and $\mathbf{z}$ are used across all methods for a given MOCOP, as summarized in Table 7.
Table 7: Reference points and ideal points for the MOCO problems.
Problem Size $\mathbf{r}$ $\mathbf{z}$
Bi-TSP 20 (20, 20) (0, 0)
50 (35, 35) (0, 0)
100 (65, 65) (0, 0)
150 (85, 85) (0, 0)
200 (115, 115) (0, 0)
Bi-CVRP 20 (30, 8) (0, 0)
50 (45, 8) (0, 0)
100 (80, 8) (0, 0)
Bi-KP 50 (5, 5) (30, 30)
100 (20, 20) (50, 50)
200 (30, 30) (75, 75)
Tri-TSP 20 (20, 20, 20) (0, 0)
50 (35, 35, 35) (0, 0)
100 (65, 65, 65) (0, 0)
D.2 Inverted Generational Distance
The Inverted Generational Distance (IGD) is a commonly used metric to assess the quality of solutions generated by multi-objective optimization methods. It captures both convergence to the Pareto front and diversity among solutions by averaging the shortest distances from each point in a reference set of Pareto-optimal solutions to the closest solution in the approximated front. Given a reference front $Q$ and a non-dominated set $\mathcal{P}$ obtained by the algorithm, IGD is computed as:
$\mathrm{IGD}=\frac{1}{|Q|}\sum_{q\in Q}\min_{p\in\mathcal{P}}\|q-p\|_{2},$ (12)
where $\|\cdot\|_{2}$ denotes the Euclidean norm. Lower IGD values indicate that the solution set $\mathcal{P}$ is closer to the true Pareto front $Q$ in terms of both convergence and diversity. In this study, the reference set $Q$ is constructed as the non-dominated front derived from the union of all heuristics (Coello Coello and Reyes Sierra 2004).
D.3 Shannon-Wiener diversity index
The Shannon–Wiener Diversity Index (SWDI) provides a quantitative measure of population diversity at a given time step, based on the distribution of individuals into clusters (Dat, Doan, and Binh 2025). In the context of heuristic search algorithms, this index reflects how evenly the population is spread across the search space. Given a set of encoded individuals $\mathcal{V}=\{\mathbf{v}_{1},\dots,\mathbf{v}_{n}\}$, each individual is assigned to a cluster using cosine similarity. Let $C_{i}$ denote the $i$-th cluster, and let $M$ be the total number of individuals in all clusters. The proportion of individuals in cluster $C_{i}$ is defined as:
$p_{i}=\frac{|C_{i}|}{M}.$
The diversity score is then computed using the Shannon entropy:
$H(X)=-\sum_{i=1}^{N}p_{i}\log(p_{i}),$
where $N$ is the total number of clusters. A higher value of $H(X)$ indicates a more uniform distribution of individuals across clusters, suggesting better exploration of the search space. In contrast, a lower value implies that individuals are concentrated in fewer clusters, which may facilitate exploitation of promising regions but also increases the risk of premature convergence. In this study, the clustering procedure and associated hyper-parameters follow the configuration described in (Dat, Doan, and Binh 2025).
D.4 Cummmulative diversity index
In the context of heuristic search, the Cumulative Diversity Index (CDI) quantifies how well the diversity, or system energy, is distributed from a centralized state to a more dispersed configuration.(Dat, Doan, and Binh 2025).
Let $A=\{\mathbf{v}_{1},\dots,\mathbf{v}_{n}\}$ be the set of all individuals in the archive, where each individual is represented by an embedding vector in a continuous space. To assess the diversity within $A$, a Minimum Spanning Tree (MST) is constructed over the set using Euclidean distances between individual vectors. The MST connects all individuals with a subset of |A|1|A|-1 edges such that the total edge length is minimized and no cycles are formed. Let $d_{i}$ denote the length of the $i$-th edge in the MST. The probability associated with each edge is computed as:
$p_{i}=\frac{d_{i}}{\sum_{j=1}^{|A|-1}d_{j}},$
and the cumulative diversity is then defined using Shannon entropy:
$H(X)=-\sum_{i=1}^{|A|-1}p_{i}\log(p_{i}),$
where $|A|$ is the archive size. Higher CDI values indicate a more distributed and diverse population, which is essential for maintaining a robust search process.
Appendix E MPaGE prompt details
E.1 Task description prompt
Our goal is to design a heuristic function for generating high-quality neighbor solutions in specific MOCOP. The task description provided in the prompt and the template of the Python code snippet is outlined below. The inputs include a solution archive and relevant problem-specific data, and the output should be a feasible neighbor solution.
Bi-TSP heuristic design task description and template program. Task Description: You are solving a Bi-objective Travelling Salesman Problem (bi-TSP), where each node has two different 2D coordinates: $(x_{1},y_{1})$ and $(x_{2},y_{2})$, representing its position in two objective spaces. The goal is to find a tour visiting each node exactly once and returning to the starting node, while minimizing two objectives simultaneously: the total tour length in each coordinate space. Given an archive of solutions, where each solution is a numpy array representing a TSP tour, and its corresponding objective is a tuple of two values (cost in each space), design a heuristic function named select_neighbor that selects one solution from the archive and applies a novel or hybrid local search operator to generate a neighbor solution from it. Please perform an intelligent random selection from among the solutions that show promising potential for further local improvement. Using a creative local search strategy that you design yourself, go beyond standard approaches to design a method that yields higher-quality solutions across multiple objectives. The function should return the new neighbor solution. Template Program: 1import numpy as np 2from typing import List, Tuple 3import random 4 5def select_neighbor( 6 archive: List[Tuple[np.ndarray, Tuple[float, float]]], 7 instance: np.ndarray, 8 distance_matrix_1: np.ndarray, 9 distance_matrix_2: np.ndarray 10) -> np.ndarray: 11 """ 12 Select a promising solution from the archive and generate a neighbor solution from it. 13 14 Args: 15 archive: List of (solution, objective) pairs. Each solution is a numpy array of node IDs. 16 Each objective is a tuple of two float values (cost in each space). 17 instance: Numpy array of shape (N, 4). Each row contains coordinates in 2D spaces: (x1, y1, x2, y2). 18 distance_matrix_1: Distance matrix in the first objective space. 19 distance_matrix_2: Distance matrix in the second objective space. 20 21 Returns: 22 A new neighbor solution (numpy array). 23 """ 24 base_solution = archive[0][0].copy() 25 new_solution = base_solution.copy() 26 new_solution[0], new_solution[1] = new_solution[1], new_solution[0] 27 28 return new_solution
Tri-TSP heuristic design task description and template program. Task Description: You are solving a Tri-objective Travelling Salesman Problem (tri-TSP), where each node has three different 2D coordinates: $(x_{1},y_{1})$, $(x_{2},y_{2})$, and $(x_{3},y_{3})$, representing its position in three objective spaces. The goal is to find a tour visiting each node exactly once and returning to the starting node, while minimizing three objectives simultaneously: the total tour length in each coordinate space. Given an archive of non-dominated solutions, where each solution is a numpy array representing a TSP tour, and its corresponding objective is a tuple of three values (cost in each space), design a heuristic function named select_neighbor that selects one solution from the archive and applies a novel or hybrid local search operator to generate a neighbor solution from it. Please perform an intelligent random selection from among the solutions that show promising potential for further local improvement. Using a creative local search strategy of your own design, specifically tailored to effectively optimize across three objectives, go beyond standard approaches to design a method that yields higher-quality solutions across multiple objectives. The function should return the new neighbor solution. Template Program: 1import numpy as np 2from typing import List, Tuple 3import random 4 5def select_neighbor( 6 archive: List[Tuple[np.ndarray, Tuple[float, float, float]]], 7 instance: np.ndarray, 8 distance_matrix_1: np.ndarray, 9 distance_matrix_2: np.ndarray, 10 distance_matrix_3: np.ndarray 11) -> np.ndarray: 12 """ 13 Select a promising solution from the archive and generate a neighbor solution from it. 14 15 Args: 16 archive: List of (solution, objective) pairs. Each solution is a numpy array of node IDs. 17 Each objective is a tuple of three float values (costs in each space). 18 instance: Numpy array of shape (N, 6). Each row contains coordinates: (x1, y1, x2, y2, x3, y3). 19 distance_matrix_1: Distance matrix in the first objective space. 20 distance_matrix_2: Distance matrix in the second objective space. 21 distance_matrix_3: Distance matrix in the third objective space. 22 23 Returns: 24 A new neighbor solution (numpy array). 25 """ 26 base_solution = archive[0][0].copy() 27 new_solution = base_solution.copy() 28 new_solution[0], new_solution[1] = new_solution[1], new_solution[0] 29 30 return new_solution
Bi-KP heuristic design task description and template program. Task Description: You are solving a Bi-objective Knapsack Problem (BI-KP), where each item has a weight and two profit values: value1 and value2. The goal is to select a subset of items such that the total weight does not exceed a given capacity, while simultaneously maximizing the total value in both objective spaces. Given an archive of non-dominated solutions, where each solution is a binary numpy array indicating item inclusion (1) or exclusion (0), and its corresponding objective is a tuple of two values (total value1, total value2), design a heuristic function named select_neighbor that selects one solution from the archive and applies a novel or hybrid local search operator to generate a neighbor solution from it. You must ensure that the generated neighbor solution remains feasible. Please perform an intelligent random selection from among the solutions that show promising potential for further local improvement. Using a creative local search strategy that you design yourself, go beyond standard approaches to develop a method that yields higher-quality solutions across multiple objectives. The function should return the new neighbor solution. Template Program: 1import numpy as np 2from typing import List, Tuple 3import random 4 5def select_neighbor( 6 archive: List[Tuple[np.ndarray, Tuple[float, float]]], 7 weight_lst: np.ndarray, 8 value1_lst: np.ndarray, 9 value2_lst: np.ndarray, 10 capacity: float 11) -> np.ndarray: 12 """ 13 Select a promising solution from the archive and generate a neighbor solution from it. 14 15 Args: 16 archive: List of (solution, objective) pairs. Each solution is a binary numpy array (0/1) of item selections. 17 Each objective is a tuple of two float values (total value1, total value2). 18 weight_lst: Numpy array of shape (N,), item weights. 19 value1_lst: Numpy array of shape (N,), item values for objective 1. 20 value2_lst: Numpy array of shape (N,), item values for objective 2. 21 capacity: Maximum allowed total weight. 22 23 Returns: 24 A new neighbor solution (numpy array). 25 """ 26 base_solution = archive[0][0].copy() 27 new_solution = base_solution.copy() 28 new_solution[0], new_solution[1] = new_solution[1], new_solution[0] 29 30 return new_solution
Bi-CVRP heuristic design task description and template program. Task Description: You are solving a Bi-objective Capacitated Vehicle Routing Problem (Bi-CVRP), where a single depot and multiple customers are located in 2D space. Each customer has a positive demand, and all vehicles in the fleet have identical capacity limits. The objective is to construct a set of routes, each starting and ending at the depot, such that: all customers are served, vehicle capacities are not exceeded on any route, two conflicting objectives are minimized: total travel distance across all routes, makespan (the length of the longest individual route). Each solution in the archive is represented as a list of NumPy arrays, where each array denotes a single route (starting and ending at depot index 0), and is paired with a tuple of two objective values (total_distance, makespan). Your task is to implement a function named select_neighbor that selects one promising solution from the archive and applies a novel or hybrid local search operator to generate a feasible neighbor solution. You must ensure that vehicle capacity constraints are respected. Please perform an intelligent random selection among solutions that show potential for local improvement. Go beyond standard approaches to develop a method that yields higher-quality solutions across both objectives. The function should return the new neighbor solution. Template Program: 1import numpy as np 2from typing import List, Tuple 3import random 4 5def select_neighbor( 6 archive: List[Tuple[np.ndarray, Tuple[float, float]]], 7 coords: np.ndarray, 8 demand: np.ndarray, 9 distance_matrix: np.ndarray, 10 capacity: float 11) -> np.ndarray: 12 """ 13 Select a promising solution from the archive and generate a neighbor solution from it. 14 Args: 15 archive: A list of tuples, where each tuple contains: 16 - solution: A list of numpy arrays, each representing a vehicle route. 17 Each route starts and ends at the depot (node index 0), e.g., [0, 3, 5, 0]. 18 - objective: A tuple of two float values (total_distance, makespan), 19 representing the two objective values of the solution. 20 21 coords: A numpy array of shape (n_nodes, 2), representing (x, y) coordinates of each node (depot + customers). 22 demand: A numpy array of shape (n_nodes,), where demand[i] is the demand of node i. The depot has demand 0. 23 distance_matrix: A numpy array of shape (n_nodes, n_nodes), where [i][j] is the Euclidean distance between node i and j. 24 capacity: A float representing the maximum capacity of each vehicle. 25 26 Returns: 27 A new neighbor solution. 28 """ 29 base_solution = archive[0][0].copy() 30 new_solution = base_solution.copy() 31 32 return new_solution
E.2 Initialization prompt
In our experiments, all initial heuristics are generated by LLMs without requiring any expert-crafted designs. The LLMs are prompted with the heuristic design task and instructed to produce new heuristics by first providing a textual description, followed by a corresponding Python implementation. This process is repeated $N$ times to obtain $N$ distinct initial heuristics.
Prompt for population initialization. You are an expert in the domain of optimization heuristics helping to design heuristics that can effectively solve optimization problems. {Task Description} 1. First, describe your new algorithm and main steps in one long, detail sentence. The description must be inside within boxed {{}. 2. Next, implement the following Python function: {Template Program} Check syntax, code carefully before returning the final function. Do not give additional explanations.
Here, Task Description and Template Program are defined in E.1.
E.3 Semantic clustering prompt
We employ the following prompt to instruct the LLM to analyze and cluster a list of heuristic code snippets based on their semantic logic. Each snippet corresponds to a heuristic generated in earlier stages and is represented as a Python function. The LLM is asked to group these snippets such that heuristics with similar behavior or design principles fall into the same cluster. This enables automated semantic analysis and categorization without manual labeling. The expected output is a JSON object, where each key corresponds to a cluster ID (as a string), and each value is a list of indices indicating the heuristics belonging to that cluster.
To improve efficiency and avoid redundant computation, we cache the clustering results for each grid cell along with its neighboring cells. These cached results can be reused in subsequent runs, substantially reducing overall processing time.
Prompt for Grouping Code Snippets You are an expert in the domain of optimization heuristics helping to design heuristics that can effectively solve optimization problems. I have {len(codes)} code snippets as follows: <Code>: ... Analyze the logic of all the given code snippets carefully. Then group the snippets into clusters where each group contains codes with similar logic. Return the result as a JSON object where the keys are the group indices and the values are lists of code indices that belong to each group. For example:
  "1": [0, 2, 4],
  "2": [1, 3],
  "3": [5]
E.4 Feedback reflection prompt
We use this prompt to guide the LLM in synthesizing a new heuristic by reflecting on a set of existing candidate heuristics. Each candidate is provided in code form, and the LLM is instructed to analyze common patterns, identify shared strengths and recurring weaknesses, and ultimately propose a single hybrid or improved strategy.
Prompt for heuristic synthesis. You are an expert in the domain of optimization heuristics helping to design heuristics that can effectively solve optimization problems. {Task Description} I have {len(indivs)} existing algorithms with their codes as follows: <Code>:  ...
Please carefully analyze all of the above algorithms. Your task is to synthesize their ideas, identify recurring patterns, and point out opportunities for improvement. Your output should be a Suggestions section, where you: Summarize key strengths shared across the implementations. Identify limitations or blind spots that appear in multiple codes. Propose hybrid or improved strategies that integrate strengths and overcome shortcomings, in a feasible running time. Output format: Suggestions: Write only one proposed hybrid or improved strategy that integrates strengths and overcomes shortcomings here. Do not include any explanations, summaries, or new algorithms outside of this section.
Here, len(indivs) is the number of selected heuristics and <Code> are their corresponding Python implementations and descripions, respectively.
E.5 Crossover prompt
MPaGE adopts search operators from EoH (Liu et al. 2024), each implemented via LLMs. Beyond this standard structure, we optionally include a reflective feedback segment, highlighted in red, which is derived from the synthesis output of E.4. This feedback provides suggestions for improvement based on analysis of previously generated heuristics. Importantly, not all operators incorporate this component: mutation operators are generated without feedback, while crossover operators may include it with a certain probability $\rho$. This probabilistic inclusion aims to balance solution quality with computational efficiency, by reducing the frequency of costly LLM inference.
E1 Operator. I have {len(indivs)} existing algorithms with their codes as follows: <Code>:  ...
Analyze the logic of all the given code snippets carefully. Then identify the two code snippets whose logic is most different from each other and create a new algorithm that is totally different in both logic and form from both of them. Here are some suggestions you can refer to:
Suggestions:
+ {suggestions} +
1. First, describe your new algorithm and main steps in one long, detailed sentence. The description must be inside within boxed {{}. 2. Next, implement the following Python function:
{Template Program}
Check syntax, code carefully before returning the final function. Do not give additional explanations.
E2 Operator. I have {len(indivs)} existing algorithms with their codes as follows: <Code>:  ...
Here are some suggestions you can refer to:
Suggestions:
+ {suggestions} +
Please help me create a new algorithm that has a totally different form from the given ones but can be motivated from them. 1. Firstly, identify the common backbone idea in the provided algorithms. 2. Secondly, based on the backbone idea, describe your new algorithm. The description must be inside within boxed {{}. 3. Thirdly, implement the following Python function:
{Template Program}
Check syntax, code carefully before returning the final function. Do not give additional explanations.
M1 Operator. I have one algorithm with its code as follows. <Code>:  ...
Please assist me in creating a new algorithm that has a different form but can be a modified version of the algorithm provided. You may focus on refining either the selection phase or the neighborhood search phase. 1. First, describe your new algorithm and main steps in one long, detailed sentence. The description must be inside within boxed {{}. 2. Next, implement the following Python function:
{Template Program}
Check syntax, code carefully before returning the final function. Do not give additional explanations.
M2 Operator. I have one algorithm with its code as follows. <Code>:  ...
Please identify the main algorithm parameters and assist me in creating a new algorithm that has a different parameter setting of the score function provided. You may focus on refining either the selection phase or the neighborhood search phase. 1. First, describe your new algorithm and main steps in one long, detailed sentence. The description must be inside within boxed {{}. 2. Next, implement the following Python function:
{Template Program}
Check syntax, code carefully before returning the final function. Do not give additional explanations.
Appendix F Clustering variations analysis
We analyze the effectiveness of different clustering methods in grouping heuristics according to their underlying logic rather than superficial code characteristics. To this end, we construct a controlled code space consisting of three ground-truth clusters, where each cluster contains multiple heuristics that implement the same functional behavior. Although all heuristics within a cluster share identical logic, they differ in implementation details such as programming constructs (e.g., use of loops vs. vectorized operations), random sampling techniques, copying mechanisms, or library functions. This design introduces syntactic and stylistic variations that challenge clustering algorithms to look beyond surface-level code differences. Illustrative examples of these heuristic groups are shown in Figures 46, where each group demonstrates the same operational intent expressed through distinct code structures. This setup allows us to evaluate whether clustering methods can correctly recover the latent functional groupings, rather than being misled by low-level syntactic variations.
We evaluate four clustering strategies:
  1. 1.
    Ours (MPaGE): Leverages LLMs to semantically interpret the code and group heuristics based on inferred logical behavior.
  2. 2.
    SWDI-based clustering: Converts code snippets into vector representations using a code embedding model and performs clustering based on similarity, following the setting in Dat, Doan, and Binh (2025).
  3. 3.
    K-Means: Clusters heuristics solely based on their performance across optimization objectives, without incorporating any code-level information. The number of clusters is fixed at 3 to align with the semantic clustering setting.
  4. 4.
    AST Similarity: Computes pairwise structural similarity between the Abstract Syntax Trees (ASTs) of the heuristics, as described in Yao, Chen, and Wang (2025). The AST similarity score ranges from 0 to 1, where 0 indicates complete structural dissimilarity and 1 signifies syntactically identical structures.
Figure 4: Group Heuristics 1: Three heuristics implementing the same logic: selecting a solution and generating a neighbor by randomly swapping two elements.
Based on Figure 7 and Figure 8, it is evident that the MPaGE method yields clustering results that most closely align with the ground-truth functional grouping. This demonstrates that LLMs, when properly utilized, are capable of understanding the semantics and intended logic of code in ways that surpass purely quantitative approaches. In contrast, the K-Means clustering method, which groups heuristics based on objective performance metrics, produces significantly less coherent clusters. This is because quantitative values, such as execution time or solution quality, do not reliably capture the underlying logic of code. Such values are often affected by stochastic behavior, system-dependent conditions, or implementation-level optimizations, leading to misleading assessments of similarity. Two heuristics implementing the same logic may yield different numerical results for reasons entirely unrelated to their semantic behavior.
The SWDI approach, which relies on embedding representations, partially captures structural similarity (e.g., correctly grouping the last three heuristics), but overall lacks consistency. Code embeddings are heavily influenced by surface-level patterns and often fail to account for the high degree of syntactic variability introduced by LLMs. As a result, heuristics with equivalent functionality may be represented as distant points in the embedding space. The AST similarity also matrix exhibits limited discriminative power. Although ASTs reflect syntactic structure, they remain sensitive to superficial differences in code, particularly in LLM-generated heuristics, where identical logic can manifest in diverse syntactic forms. As observed in Figure 7e, the pairwise similarity scores between heuristic pairs fail to highlight or emphasize any significant structural grouping. Similarity values tend to be uniformly distributed and do not correspond well to the functional relationships among heuristics.
Figure 5: Group Heuristics 2: Three heuristics implementing the same logic: selecting a solution and reversing a randomly chosen segment $[i:j+1]$.
Figure 6: Group Heuristics 3: Four heuristics implementing the same logic: Sequentially select unvisited nodes such that the next node in the closet with the smallest sum of total distance.
(a) Ground Truth
(b) MPaGE
(c) SWDI Approach
(d) K-Means
(e) AST Similarity
Figure 7: Clustering consistency and similarity matrices across various methods. Each heatmap represents pairwise relationships between code heuristics. In subfigures (a)–(d), a value of 1 indicates that two heuristics are assigned to the same cluster, while 0 indicates they are assigned to different clusters. (a) Ground-truth grouping by functional logic. (b) Clustering result from our method (MPaGE). (c) Clustering based on SWDI. (d) K-Means clustering based on objective performance. (e) Pairwise AST similarity scores between heuristics, where higher values indicate greater syntactic similarity between the corresponding Abstract Syntax Trees.
Table 8: Performance comparison on Bi-TSP20 and Tri-TSP20 instances.
Method Bi-TSP20 Tri-TSP20
HV $\uparrow$ IGD $\downarrow$ HV $\uparrow$ IGD $\downarrow$
SWDI Cluster 0.911 0.024 0.888 0.098
K-Means 0.876 0.030 0.815 0.184
AST Similarity 0.745 0.184 0.757 0.203
MPaGE (Ours) 0.921 0.013 0.892 0.102
To evaluate the effectiveness of different strategies in guiding parent selection for heuristic synthesis, we conduct experiments on two benchmark problems: Bi-TSP20 and Tri-TSP20, as shown in Table 8. For the SWDI Clustering and K-Means baselines, parent selection is performed based on cluster membership, similar to MPaGE. In contrast, the AST Similarity method samples parent pairs based on their pairwise structural similarity scores.
As shown in Table 8, MPaGE consistently outperforms all baselines across both benchmark instances. On Bi-TSP20, MPaGE achieves the highest HV (0.921) and the lowest IGD (0.013), indicating superior convergence and solution diversity. A similar pattern is observed on Tri-TSP20, where MPaGE attains the best HV (0.892) and a competitive IGD (0.102), closely approaching the best IGD achieved by SWDI Clustering (0.098). These results suggest that relying solely on AST similarity offers limited utility for parent selection, particularly when dealing with heuristics that involve complex or subtle logic structures.
(a) Ground Truth Clusters
(b) MPaGE Clusters
(c) SWDI Clusters
(d) K-Means
Figure 8: Visualization of heuristic clustering results produced by different methods. (a) Ground truth clusters based on manual semantic annotation. (b) MPaGE clusters, which are generated using LLM-driven semantic interpretation of code logic. (c) SWDI clusters, derived from code embeddings using an embedding model. (d) K-Means clusters, obtained by grouping heuristics based solely on performance metrics. MPaGE most closely approximates the ground truth structure, while other methods exhibit greater semantic drift, especially in cases with subtle logic differences.
Appendix G Designed Heuristics
This section presents the best-performing heuristics generated by MPaGE in terms of hypervolume and runtime across all benchmark instances.
Bi-TSP
# Best hypervolume
# score = [0.584, 0.505]
1def select_neighbor(archive: List[Tuple[np.ndarray, Tuple[float, float]]],
2 instance: np.ndarray,
3 distance_matrix_1: np.ndarray,
4 distance_matrix_2: np.ndarray) -> np.ndarray:
5 # Compute selection probabilities based on inverse objective values
6 total_objective_value = sum(1 / (obj[0] + 1e-9) + 1 / (obj[1] + 1e-9)
7 for _, obj in archive)
8 probabilities = [(1 / (obj[0] + 1e-9) + 1 / (obj[1] + 1e-9)) / total_objective_value
9 for _, obj in archive]
10
11 # Select a solution from the archive based on computed probabilities
12 selected_index = np.random.choice(len(archive), p=probabilities)
13 selected_solution = archive[selected_index][0].copy()
14
15 # Generate neighbor using segment reversal
16 n = len(selected_solution)
17 i, j = sorted(random.sample(range(n), 2)) # Choose two distinct indices
18 new_solution = np.concatenate((selected_solution[:i],
19 selected_solution[i:j+1][::-1],
20 selected_solution[j+1:]))
21
22 return new_solution
# Best runningtime
# score = [0.395, 0.121]
1def select_neighbor(archive: List[Tuple[np.ndarray, Tuple[float, float]]],
2 instance: np.ndarray,
3 distance_matrix_1: np.ndarray,
4 distance_matrix_2: np.ndarray) -> np.ndarray:
5 # Compute selection probabilities based on inverse objective values
6 total_score = sum(1 / (obj[0] + 1e-9) + 1 / (obj[1] + 1e-9)
7 for _, obj in archive)
8 probabilities = [(1 / (obj[0] + 1e-9) + 1 / (obj[1] + 1e-9)) / total_score
9 for _, obj in archive]
10
11 # Select a solution from the archive based on computed probabilities
12 selected_index = np.random.choice(len(archive), p=probabilities)
13 selected_solution = archive[selected_index][0].copy()
14
15 # Generate neighbor using node repositioning
16 n = len(selected_solution)
17 neighbor_solution = selected_solution.copy()
18
19 # Select two distinct nodes
20 idx1, idx2 = random.sample(range(n), 2)
21
22 # Heuristic: move node1 right after, node2 right before
23 new_position1 = (idx1 + 1) % n
24 new_position2 = (idx2 - 1) % n
25
26 # Apply the swap
27 neighbor_solution[new_position1], neighbor_solution[new_position2] = neighbor_solution[idx1], neighbor_solution[idx2]
28
29 return neighbor_solution
Tri-TSP
# Best hypervolume
# score = [0.359, 0.893]
1def select_neighbor(archive: List[Tuple[np.ndarray, Tuple[float, float, float]]],
2 instance: np.ndarray,
3 distance_matrix_1: np.ndarray,
4 distance_matrix_2: np.ndarray,
5 distance_matrix_3: np.ndarray) -> np.ndarray:
6 # Select a promising solution based on multi-objective performance
7 archive_weights = [1 / (1 + sum(obj)) for _, obj in archive]
8 weighted_archive = random.choices(archive, weights=archive_weights, k=1)[0]
9 base_solution = weighted_archive[0].copy()
10
11 new_solution = base_solution.copy()
12 N = len(new_solution)
13
14 # Compute performance score and set perturbation factor
15 first_objective, second_objective, third_objective = weighted_archive[1]
16 avg_objective = (first_objective + second_objective + third_objective) / 3
17 perturbation_factor = max(0.1, 0.5 + (0.5 - avg_objective))
18
19 # Diversify neighborhood: Swap, Reverse, or Shift
20 mutation_type = random.choices([’swap’, ’reverse’, ’shift’],
21 weights=[0.5 * perturbation_factor,
22 0.3 * (1 - perturbation_factor),
23 0.2],
24 k=1)[0]
25
26 if mutation_type == ’swap’:
27 idx1, idx2 = random.sample(range(1, N - 1), 2)
28 new_solution[idx1], new_solution[idx2] = new_solution[idx2], new_solution[idx1]
29 elif mutation_type == ’reverse’:
30 start_idx = random.randint(1, N - 2)
31 end_idx = random.randint(start_idx + 1, N - 1)
32 new_solution[start_idx:end_idx + 1] = new_solution[start_idx:end_idx + 1][::-1]
33 elif mutation_type == ’shift’:
34 shift_idx = random.randint(1, N - 2)
35 new_solution[shift_idx], new_solution[shift_idx + 1] = new_solution[shift_idx + 1], new_solution[shift_idx]
36
37 # Adaptive large perturbation for good solutions
38 if avg_objective < 0.5:
39 perturb_indices = random.sample(range(1, N - 1), k=min(3, N - 2))
40 random.shuffle(perturb_indices)
41 for i in range(len(perturb_indices) - 1):
42 new_solution[perturb_indices[i]], new_solution[perturb_indices[i + 1]] = (
43 new_solution[perturb_indices[i + 1]], new_solution[perturb_indices[i]]
44 )
45
46 # Additional large shuffle with some probability
47 if random.random() < 0.3:
48 additional_indices = random.sample(range(1, N - 1), k=3)
49 random.shuffle(additional_indices)
50 new_solution[additional_indices[0]], new_solution[additional_indices[1]], new_solution[additional_indices[2]] = (
51 new_solution[additional_indices[1]], new_solution[additional_indices[2]], new_solution[additional_indices[0]]
52 )
53
54 return new_solution
# Best running time
# score = [0.335, 0.476]
1def select_neighbor(archive: List[Tuple[np.ndarray, Tuple[float, float, float]]],
2 instance: np.ndarray,
3 distance_matrix_1: np.ndarray,
4 distance_matrix_2: np.ndarray,
5 distance_matrix_3: np.ndarray) -> np.ndarray:
6 # Select a solution with good objective sum
7 selected_solution, _ = random.choices(
8 archive,
9 weights=[1 / (obj[0] + obj[1] + obj[2]) for _, obj in archive],
10 k=1
11 )[0]
12
13 # Apply a simple swap
14 neighbor_solution = selected_solution.copy()
15 n = len(neighbor_solution)
16 i, j = random.sample(range(n), 2)
17 neighbor_solution[i], neighbor_solution[j] = neighbor_solution[j], neighbor_solution[i]
18
19 return neighbor_solution
Bi-KP
# Best hypervolume
# score = [0.348, 0.130]
1def select_neighbor(archive: List[Tuple[np.ndarray, Tuple[float, float]]],
2 weight_lst: np.ndarray,
3 value1_lst: np.ndarray,
4 value2_lst: np.ndarray,
5 capacity: float) -> np.ndarray:
6 # Select a solution with high objective sum
7 selected_pair = random.choices(
8 archive,
9 weights=[sol[1][0] + sol[1][1] for sol in archive],
10 k=1
11 )[0]
12
13 base_solution = selected_pair[0].copy()
14 new_solution = base_solution.copy()
15 current_weight = np.dot(new_solution, weight_lst)
16
17 selected_indices = np.where(new_solution == 1)[0]
18 unselected_indices = np.where(new_solution == 0)[0]
19
20 # Prioritize swapping items with high profit-to-weight ratio
21 if selected_indices.size > 0 and unselected_indices.size > 0:
22 profit_to_weight = (value1_lst[selected_indices] + value2_lst[selected_indices]) / weight_lst[selected_indices]
23 sorted_selected_indices = selected_indices[np.argsort(profit_to_weight)[::-1]]
24
25 for _ in range(5): # Try multiple swaps for improvement
26 selected_idx = random.choice(sorted_selected_indices)
27 unselected_idx = random.choice(unselected_indices)
28
29 new_solution[selected_idx] = 0
30 new_solution[unselected_idx] = 1
31
32 if np.dot(new_solution, weight_lst) <= capacity:
33 return new_solution # Accept valid neighbor
34
35 # Revert swap if constraint violated
36 new_solution[selected_idx] = 1
37 new_solution[unselected_idx] = 0
38
39 # Fallback: perturb multiple items
40 num_toggles = random.randint(2, 4)
41 for _ in range(num_toggles):
42 idx = random.randint(0, len(base_solution) - 1)
43 if new_solution[idx] == 1:
44 new_solution[idx] = 0
45 else:
46 if current_weight + weight_lst[idx] <= capacity:
47 new_solution[idx] = 1
48 current_weight += weight_lst[idx]
49
50 # Final validation: enforce capacity
51 while np.dot(new_solution, weight_lst) > capacity:
52 deselect_idx = random.choice(np.where(new_solution == 1)[0])
53 new_solution[deselect_idx] = 0
54
55 return new_solution
# Best running time
# score = [0.319, 0.074]
1def select_neighbor(archive: List[Tuple[np.ndarray, Tuple[float, float]]],
2 weight_lst: np.ndarray,
3 value1_lst: np.ndarray,
4 value2_lst: np.ndarray,
5 capacity: float) -> np.ndarray:
6 # Randomly sample a base solution
7 selected_solution, _ = random.choice(archive)
8 neighbor_solution = selected_solution.copy()
9 num_items = len(weight_lst)
10
11 # Flip 1 to 3 random items with feasibility check
12 for _ in range(random.randint(1, 3)):
13 item_index = random.randint(0, num_items - 1)
14 neighbor_solution[item_index] = 1 - neighbor_solution[item_index]
15
16 # Undo flip if capacity exceeded
17 while np.dot(neighbor_solution, weight_lst) > capacity:
18 neighbor_solution[item_index] = 1 - neighbor_solution[item_index]
19 item_index = random.randint(0, num_items - 1)
20 neighbor_solution[item_index] = 1 - neighbor_solution[item_index]
21
22 return neighbor_solution
Bi-CVRP
# Best hypervolume
# score = [0.417, 0.158]
1def select_neighbor(archive: List[Tuple[np.ndarray, Tuple[float, float]]],
2 coords: np.ndarray,
3 demand: np.ndarray,
4 distance_matrix: np.ndarray,
5 capacity: float) -> np.ndarray:
6 # Select a solution with best makespan (2nd objective)
7 best_solution, _ = min(archive, key=lambda x: x[1][1])
8 neighbor_solution = [np.copy(route) for route in best_solution]
9
10 # Choose two distinct routes
11 route_from_index = np.random.choice(len(neighbor_solution))
12 route_to_index = np.random.choice([i for i in range(len(neighbor_solution)) if i != route_from_index])
13
14 if len(neighbor_solution[route_from_index]) > 2 and len(neighbor_solution[route_to_index]) > 2:
15 # Select customers (excluding depot)
16 customer_from_index = np.random.randint(1, len(neighbor_solution[route_from_index]) - 1)
17 customer_to_index = np.random.randint(1, len(neighbor_solution[route_to_index]) - 1)
18
19 customer_from = neighbor_solution[route_from_index][customer_from_index]
20 customer_to = neighbor_solution[route_to_index][customer_to_index]
21
22 # Swap the customers
23 neighbor_solution[route_from_index][customer_from_index] = customer_to
24 neighbor_solution[route_to_index][customer_to_index] = customer_from
25
26 # Validate route demands
27 demand_from = np.sum(demand[neighbor_solution[route_from_index]])
28 demand_to = np.sum(demand[neighbor_solution[route_to_index]])
29
30 if demand_from > capacity or demand_to > capacity:
31 # Revert if infeasible
32 neighbor_solution[route_from_index][customer_from_index] = customer_from
33 neighbor_solution[route_to_index][customer_to_index] = customer_to
34
35 return neighbor_solution
# Best running time
# score = [0.146, 0.033]
1def select_neighbor(archive: List[Tuple[np.ndarray, Tuple[float, float]]],
2 coords: np.ndarray,
3 demand: np.ndarray,
4 distance_matrix: np.ndarray,
5 capacity: float) -> np.ndarray:
6 # Select route with smallest makespan
7 min_makespan_solution = min(archive, key=lambda x: x[1][1])
8 routes = min_makespan_solution[0]
9
10 # Choose two distinct routes
11 route1_index, route2_index = np.random.choice(len(routes), 2, replace=False)
12 route1 = routes[route1_index]
13 route2 = routes[route2_index]
14
15 if len(route1) > 2 and len(route2) > 2:
16 # Choose customers (excluding depot)
17 customer1_index = np.random.randint(1, len(route1) - 1)
18 customer2_index = np.random.randint(1, len(route2) - 1)
19
20 customer1 = route1[customer1_index]
21 customer2 = route2[customer2_index]
22
23 # Attempt to swap
24 new_route1 = route1.copy()
25 new_route2 = route2.copy()
26 new_route1[customer1_index], new_route2[customer2_index] = customer2, customer1
27
28 # Check if feasible
29 demand1 = np.sum(demand[new_route1[1:-1]])
30 demand2 = np.sum(demand[new_route2[1:-1]])
31
32 if demand1 <= capacity and demand2 <= capacity:
33 routes[route1_index] = new_route1
34 routes[route2_index] = new_route2
35 else:
36 # Fallback: remove and insert
37 idx_remove = np.random.randint(1, len(route1) - 1)
38 customer = route1[idx_remove]
39 new_route1 = np.delete(route1, idx_remove)
40
41 if np.sum(demand[new_route1[1:-1]]) + demand[customer] <= capacity:
42 routes[route1_index] = new_route1
43 routes[route2_index] = np.insert(route2, -1, customer)
44
45 return routes
Appendix H Comparison to Neural Combinatorial Optimization
Table 9: Performance comparison across all benchmarks against Neural Combinatorial Optimization methods.
Method Bi-TSP20 Bi-TSP50 Bi-TSP100 Tri-TSP20 Tri-TSP50 Tri-TSP100
HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$
PMOCO 0.632 0.31 3.739 0.637 0.93 5.405 0.699 1.82 7.952 0.470 1.67 4.448 0.436 30.68 5.596 0.490 4.10 9.296
NHDE-P 0.630 0.63 3.373 0.510 20.68 9.250 0.703 1.26 11.907 0.474 0.83 14.580 0.629 0.00 25.123 0.506 0.97 65.506
NHDE-M 0.634 0.00 93.421 0.643 0.00 157.843 0.712 0.00 328.562 0.476 0.41 1195.237 0.447 28.93 1498.893 0.511 0.00 3612.458
MPaGE (best) 0.629 0.78 4.429 0.542 15.70 6.950 0.442 37.92 12.719 0.478 0.00 8.112 0.220 65.02 14.747 0.109 78.66 26.394
Method Bi-KP50 Bi-KP100 Bi-KP200 Bi-CVRP20 Bi-CVRP50 Bi-CVRP100
HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$ HV $\uparrow$ Gap Time $\downarrow$
PMOCO 0.378 0.00 5.415 0.441 9.25 8.597 0.350 5.14 11.060 0.407 28.34 5.221 0.413 9.03 9.684 0.329 22.03 14.759
NHDE-P 0.371 1.85 4.681 0.431 11.31 8.224 0.359 2.71 14.092 0.411 27.64 3.509 0.432 4.84 5.981 0.380 9.95 10.747
NHDE-M 0.353 6.61 283.514 0.453 6.79 502.367 0.369 0.00 897.214 0.421 25.88 213.678 0.441 2.86 299.112 0.396 6.16 655.309
MPaGE (best) 0.359 5.02 2.647 0.486 0.00 1.859 0.197 46.61 2.118 0.568 0.00 0.072 0.454 0.00 0.191 0.422 0.00 0.359
[…]